1. A scientist measures the standard enthalpy change for the following reaction to be -380.6 kJ :
P4O10(s) + 6 H2O(l)4H3PO4(aq)
Based on this value and the standard enthalpies of formation for the other substances, the standard enthalpy of formation of H2O(l) is ___kJ/mol.

2. A scientist measures the standard enthalpy change for the following reaction to be -578.5 kJ :
H2CO(g) + O2(g)CO2(g) + H2O(l)
Based on this value and the standard enthalpies of formation for the other substances, the standard enthalpy of formation of CO2(g) is ___kJ/mol.
1
Expert's answer
2019-10-31T07:29:29-0400

1)

"P_4O_{10}(s) + 6H_2O (l) \rightarrow 4H_3PO_4(aq)"

"\Delta H_{rxn} = \sum \Delta H^0_f(products) - \sum \Delta H^0_f(reactants)"

"\Delta H^0_f(P_4O_{10}(s)) = -2984.0 \frac{kJ}{mol}"


"\Delta H^0_f (H_3PO_4(aq)) = -1266.9 \frac{kJ}{mol}"



"-380.6 = 4\times(-1266.9) - [1\times(-2984.0)+6\times \Delta H^0_f(H_2O(l))]"

"\Delta H^0_f(H_2O(l)) = -283.8 \frac{kJ}{mol}"

2)

"H_2CO(g) +O_2(g)\rightarrow CO_2(g) + H_2O(l)"


"\Delta H_{rxn} = \sum \Delta H ^0_f(products) - \sum \Delta H^0_f(reactants)"

"\Delta H^0_f(H_2O(l)) = -285.8 \frac{kJ}{mol}"


"\Delta H^0_f(H_2CO(g)) = -116 \frac{kJ}{mol}"



"-578.5 = [1\times\Delta H^0_f(CO_2(g))+1\times(-285.8)]-[1\times(-116)+1\times0]"


"\Delta H^0_f (CO_2(g)) = -408.9 \frac{kJ}{mol}"